OpenStudy (anonymous):

A basket contains 11 pieces of fruit: 4 apples, 5 oranges, and 2 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?

OpenStudy (anonymous):

Well Jonas has a 5/11 If he doesn't replace his piece of fruit in the basket then Beth has a 4/10 If he does replace his fruit she has a 4/11

Directrix (directrix):

Probability is the number of desired outcomes divided by the number of possible outcomes. Of the 11 pieces of fruit, there are 5 oranges. Once Jonas takes a piece of fruit, 10 pieces remain. The probability of the compound event that Jonas selects an Orange AND Angie selects an apple) = 5/11 * 4/10 = 20/110 which simplified to ? @anaivillegas